OpenStudy (anonymous):

Find the area of the region enclosed by one loop of the curve. r=2sin5θ

OpenStudy (anonymous):

A loop looks like goes from 0 to pi/5 so integrate 1/2 (3 sin 5 theta) over that interval should give you the area.

OpenStudy (anonymous):

can you help me solve that step by step?

OpenStudy (anonymous):

Actually, I made a couple booboos there now I am looking at it again Should be 1/2 (2 sin 5 theta)^2 ie just 2 sin^2 (5 theta)

OpenStudy (anonymous):

okay i got that far.

OpenStudy (anonymous):

Area of one loop = pi/5 ?

OpenStudy (anonymous):

Maybe you can use double angle formula to get an easier integral....

OpenStudy (anonymous):

I haven't actually worked it out myself....

OpenStudy (anonymous):

sin^2 5theta = (half theta - sin 10 theta)/20 maybe...

OpenStudy (anonymous):

sin switches to cos in the double angle formula

OpenStudy (experimentx):

\[ 1/2 \int_{\theta_0}^{\theta_1} r^2 d\theta\]

OpenStudy (anonymous):

is the answer pi/5 ?`

OpenStudy (experimentx):

what do you mean by one loop http://www.wolframalpha.com/input/?i=polar+plot+r%3D2sin%285theta%29 ??

OpenStudy (anonymous):

one leaf/petal of the flower shape

OpenStudy (experimentx):

\[ \theta_0 = 0 \] \[ \theta_1 = \pi/5 \]

OpenStudy (anonymous):

yes those are the theta bounds for the integral

OpenStudy (anonymous):

okay thank you!!!

OpenStudy (experimentx):

yw